OpenStudy (anonymous):

find the solution for: log(3x-1)+log2=log4+log(x+2)

OpenStudy (slaaibak):

Use this: log a + log b = log ab

OpenStudy (rulnick):

log(3x-1)-log(x+2) = log4 - log2 log((3x-1)/(x+2)) = log(4/2) (3x-1)/(x+2) = 2 3x-1=2x+4 x=5

OpenStudy (campbell_st):

simplify the logs 1st log(3x-1) + log 2 = log(2(3x -1) or log(6x - 2) log 4 + log( x + 2) = log(4(x+2) = log(4x + 8) now equate both sides and solve for x 6x - 2 = 4x + 8 2x = 10 x = 5

OpenStudy (anonymous):

Thanks:) would you solve log(4x-1)=log(x+1)+log2 the same way

OpenStudy (campbell_st):

log(x+1) + log 2 = log(2(x+1) or log(2x + 2) so log(4x -1) = log(2x +2) equate and solve for x 4x -1 = 2x + 2 2x = 3 x = 3/2